Can you help me solve this word problem?
The sum of the present ages of a mother and her daughter, Jane, is 43 years. In 14 years' time, Jane will be half her mother's present age. Form an equation and solve it to find Jane's present age.

Answer by mananth(16946)

let Jane's age be x
mother 's age =y
x+y =43 ................(1)
after 14 years
Jane's age = x+14

1/2 * (y)=(x+14)
multiply by 2
y=2(x+14)
y=2x+28
y-2x=28...................2
solve equation (1) & (2)
x+y =43 multiply by 2
2x+2y=86
y-2x=28
Add
3y=114
/3
y=38 Mother's age
Plug y in any equation
daughter's age = 5 years